OpenStudy (anonymous):

A die is rolled 20 times and the number of twos that come up is tallied. Find the probability of getting exactly four twos. A. 0.202 B. 0.075 C. 0.101 D. 0.083

OpenStudy (kinggeorge):

I believe this requires use of the binomial distribution with \(n=20, k=4, p=1/6\). In that case, your probability is given by the formula \[\binom{20}{4} \cdot \left({1 \over 6}\right)^4 \left({5 \over 6}\right)^{20-4}\]

OpenStudy (kinggeorge):

If you work this out, you get a probability of approximately 0.202

OpenStudy (anonymous):

thanx

OpenStudy (kinggeorge):

You're welcome.